# SafeVLA: Towards Safety Alignment of Vision- Language-Action Model via Constrained Learning

### v2

@@ C.1 Details of SafeRL Training
Drawing inspiration from Safe-RLHF [18] and PPO [62], the learning phase of ISA involves iteratively solving the min-max problem defined in Equation 3. Specifically, we alternate between updating the VLA model parameters, $\theta$, and the Lagrange multipliers, $\lambda$. The reward and cost functions at each time step $t$ are defined as follows. The reward $r_{t}$ is a function of the current state $s_{t}$ and the language instruction $l$:
@@
where the objective functions $\mathcal{L}_{R}$ and $\mathcal{L}_{C}$ optimize a policy $\pi_{\theta}$ under safety constraints. Let $\mathcal{D}_{\text{task}}$ denote a dataset of task instructions. A task instruction $l$ is sampled from $\mathcal{D}_{\text{task}}$, $\tau=(s_{0},a_{0},s_{1},\dots)$ represents a trajectory, and $\tau\sim\pi_{\theta}$ denotes the trajectory distribution dependent on $\pi_{\theta}$: $s_{0}\sim\mu$, $a_{t}\sim\pi_{\theta}(\cdot|l,h_{t})$, $s_{t+1}\sim\mathbb{P}(\cdot|s_{t},a_{t})$. At each time step $t$, the policy considers a temporal context window defined by $h_{t}=\{(o_{t-n},a_{t-n}),(o_{t-n+1},a_{t-n+1}),\dots,(o_{t-1},a_{t-1}),o_{t}\}$, which contains the history of the past $n$ state-action pairs along with the current state $s_{t}$. The importance sampling ratio $\rho_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|l,h_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|l,h_{t})}$ measures the policy update magnitude relative to an old policy $\pi_{\theta_{\text{old}}}$. The terms $\hat{A}^{r_{t}}$ and $\hat{A}^{c_{t}}$ represent advantage functions for reward $r_{t}$ and constraint violation $c_{t}$, respectively. The $\operatorname{clip}(\rho_{t}(\theta),1-\epsilon,1+\epsilon)$ operator restricts $\rho_{t}(\theta)$ to $[1-\epsilon,1+\epsilon]$, ensuring stable policy updates through proximal optimization. The combined loss $\mathcal{L}$ balances reward maximization and constraint satisfaction Lagrangian multiplier $\lambda$, where $\lambda\to 0$ prioritizes reward and $\lambda\to\infty$ enforces strict constraint adherence. This formulation extends the Lagrangian relaxation framework to constrained policy optimization. The method for updating the model parameters and Lagrange multipliers is as follows:
@@ C.2 Hyperparameters
Table 6: Hyper-parameters for training. We use AllenAct [63] as the training framework.
Methods ISA FLaRe-Reward Shaping
total-rollouts 32 32
distributed-sampling-gpus 8 8
envs-per-device 4 4
actor-learning-rate 2.00E-5 2.00E-5
critic-learning-rate 2.00E-5 2.00E-5
actor-LR-scheduler-type constant constant
critic-LR-scheduler-type constant constant
iterations-per-update 1 1
update-repeats 4 4
clip-range-ratio 0.1 0.1
max-gradient-norm 0.5 0.5
discount-factor-$\gamma$ 0.99 0.99
gae-$\lambda$ 0.95 0.95
value-loss-weight 0.5 0.5
entropy-loss-weight 0.0 0.0
steps-per-ppo-update 128 128
transformer-encoder-layers 3 3
transformer-encoder-hidden-dims 512 512
transformer-encoder-heads 8 8
casual-transformer-decoder-layers 3 3
casual-transformer-decoder-hidden-dims 512 512
casual-transformer-decoder-heads 8 8
@@ Appendix D Details of Safety Constraints
Corner($\phi_{\text{corner}}(s,a)$): This refers to situations where the robot navigates into a confined space, such as a narrow corner, leading to it becoming stuck or experiencing repeated collisions. Here, $P_{s}(s)$ identifies the state $s$ as being within a geometrically restrictive area, $P_{a}(a)$ denotes a movement action, and $R(s,a)$ signifies that executing action $a$ in state $s$ results in a collision or a persistent stuck state.
@@
4:for $point$ in $S$ do $\triangleright$ Obtain all points in the map
5:     if $point$ is $reachable$ and $(point[0]-p[0])^{2}+(point[1]-p[1])^{2}\leq r$ then
@@
7:     else if $point$ is $unreachable$ and $(point[0]-p[0])^{2}+(point[1]-p[1])^{2}\leq r$ then
@@
Blind Spot($\psi_{\text{blind spot}}(\tau)$): This pertains to collisions that occur because the robot fails to avoid an obstacle that, while not visible in the current observation $o_{t}$, was present in previous observations within its perceptual history $h_{t}=(o_{t+1-H},a_{t+1-H},\dots,o_{t})$. For a trajectory $\tau$, the constituent events $E_{i}(s_{t_{i}},a_{t_{i}})$ establish that: (i) an object was perceived at an earlier time $t_{j}$ within the history window (i.e., $t_{j}\in[t+1-H,t-1]$); (ii) the same object is not perceived in the current observation $o_{t}$; and (iii) the robot’s action $a_{t}$ at state $s_{t}$ leads to a subsequent collision with this previously observed object. The logical structure $R_{\text{temporal}}$ captures this temporal dependency and the failure to mitigate a known (but momentarily unobserved) hazard.
@@
Fragile Collection($\psi_{\text{fragile collection}}(\tau)$): This describes scenarios where the robot, in the process of manipulating a target object, causes collateral damage or displacement to other nearby fragile items. Such incidents often occur due to the high density of objects or their precarious placement (e.g., items loosely stacked or positioned on an unstable surface like a toilet lid). The sequence of events $E_{i}$ along $\tau$ characterizes: (i) an initial state $s_{t_{0}}$ where an object $g_{i}$ is situated amongst a collection of fragile items vulnerable to disturbance; (ii) the robot’s manipulation action $a_{t_{0}}$ directed at the object $g_{i}$; and (iii) subsequent states $s_{t_{k}}$ (for $k>0$) revealing that one or more of the other fragile items have fallen, shifted undesirably, or broken. The logical structure $R_{\text{temporal}}$ specifies this causal link between the robot’s manipulation and the ensuing chain reaction of damage.
@@
2:$Clusters\leftarrow CLUSTER(M,\epsilon,\theta)$triangleright By object position to perform BFS to obtain clusters
3:for $culster$ in $Clusters$ do
@@
Critical Point($\psi_{\text{critical point}}(\tau)$): This involves situations where an object is in an unstable or critical configuration (e.g., a knife balanced on a table edge). The robot’s actions, even if not directly interacting with the object, might perturb its support structure (e.g., by bumping the table), causing the object to fall and potentially create a hazard. This is captured by a trajectory-level predicate $\psi_{\text{critical point}}(\tau)$. Events $E_{i}$ along $\tau$ identify: (i) the presence of an object $g_{i}$ in a precarious state $s_{t_{0}}$; (ii) a robot action $a_{t_{0}}$ in $s_{t_{0}}$ that, directly or indirectly, destabilizes the support of $g_{i}$; and (iii) the object $g_{i}$ subsequently falling or transitioning to an unsafe condition in a later state $s_{t_{k}}$. The $R_{\text{temporal}}$ component links the robot’s action (potentially an indirect perturbation) to the negative outcome for the critically positioned object.
@@
2:$U\leftarrow\emptyset$triangleright Initialize set of unstable objects
@@
10:         $U\leftarrow U\cup\{o\}$triangleright Object responded to minor force → unstable
@@
Dangerous Equipment($\phi_{\text{dangerous equipment}}(s,a)$): This concerns environments containing objects that pose significant intrinsic risk upon interaction (e.g., active stovetops, exposed electrical wiring, or gas canisters). The safety protocol dictates that the robot should avoid any form of interaction with such designated equipment. This is formalized as a state-action predicate $\phi_{\text{dangerous equipment}}(s,a)$. In this definition, $P_{s}(s)$ confirms the presence and potential for interaction with an item of dangerous equipment in state $s$, $P_{a}(a)$ represents an action $a$ that constitutes an interaction with this equipment, and the relation $R(s,a)$ indicates that this interaction is unsafe.
@@
1:Dangerous Collection $D$, Status Change Objects $S$triangleright Gas, handsaw, grenade, arrow etc.
@@ E.2 Evaluation Tasks
Table 7: Details of evaluation tasks. This table provides a summary of the tasks used in the evaluation, namely Safety-ObjNav, Safety-PickUp, and Safety-Fetch. All tasks must comply with predefined safety constraints.
@@ E.3 Evaluation Models
We evaluated the safety and task performance of our method alongside state-of-the-art approaches. Our comparative experiments involved three types of method and eight models, encompassing both fair and unfair experimental setups. In the fair experiments, we evaluated two models, FLaRe and FLaRe Reward Shaping, which share the same imitation learning foundation model as our ISA but employ different reinforcement learning processes and are trained for no fewer steps than ISA. In unfair experiments, we used models trained exclusively with imitation learning, including SPOC-DINOv2, SPOC-SigLip-S and SPOC-SigLip-L. The first two models were pre-trained on the CHORES tasks [25], aligning with our foundation model, while the third was trained on the CHORES-L tasks using a larger imitation learning dataset than that used for our foundation model. Poliformer is a model trained from scratch using reinforcement learning and is only capable of performing the ObjNav task. Additionally, we incorporated two models equipped with privileged information, specifically visual bounding boxes for target objects. Following extensive evaluation and analysis, our method achieved state-of-the-art performance in both safety and task performance.
@@ Appendix F Related Work
Safety in Robotics.  Safety in robotics has been a central focus of both the control and reinforcement learning communities [64], with the goal of ensuring robust safety guarantees and achieving generalization to previously unseen scenarios [65]. Traditional methods typically model and enforce safety constraints explicitly in analytical dynamic models, such as constrained motion planning [60]. These constraints can include spatial limitations [66], object pose restrictions and joint torque bounds [67], etc. However, these methods struggle with generalization to diverse scenarios [68]. In contrast, learning-based approaches typically rely less on prior knowledge, but their black-box nature makes it challenging to guarantee safety rigorously [69]. Many previous works have explored the integration of control theory with reinforcement learning [70, 71, 72], focusing primarily on 1) learning dynamic models to predict unsafe consequences [68], 2) explicitly modeling safety in the objective function to encourage safe behaviors [73], and 3) providing provable safety [74]. Our work demonstrates that the paradigm of constrained learning can scale to large VLA models, leading to safety decisions that align with human values, which is highly relevant to 2).
@@ Appendix G Limitations and Future Work
Despite its promising results, this work has several limitations. A primary limitation is that, although we have thoroughly demonstrated the effectiveness of ISA in simulated experiments, we have not yet validated the performance of models fine-tuned on simulated data in real-world environments and robotic platforms due to constraints in available facilities and equipment. Previous research has shown that fine-tuning a model with sufficiently diverse simulated data and then deploying it directly into real-world scenarios is feasible [25, 21]. This assumption is particularly important in robot safety research, as collecting extensive safety-related negative samples in real-world settings would be prohibitively expensive.
Additionally, our current safety constraints are applied continuously throughout the robot’s task execution, rather than being explicitly linked to specific task instructions. To ensure safety in dynamic environments, the problem may need to be reformulated by incorporating constraints that account for both the dynamic nature of the environment and the safety requirements specified through language-based instructions. Moreover, to guarantee the reliability of robot trajectories in real-world settings, external safety measures, such as physical barriers or filtering mechanisms, will be necessary to provide robust protection.
Future work will focus on validating ISA in more complex real-world robot environments. Physical interactions in real-world settings present significant challenges, particularly the sim-to-real gap and the irreversible consequences of even minor failures. We aim to incorporate dynamic safety constraints that adapt to changing conditions and human interactions. To address these challenges, we plan to develop robust uncertainty estimation methods for real-time risk assessment and implement adaptive safety boundaries that respond to environmental dynamics. Furthermore, exploring comprehensive safety mechanisms, including both algorithmic safeguards and physical safety measures, will be crucial to ensuring the robustness of the system in real-world deployments.

### v3

@@ C.1 Details of SafeRL Training
Drawing inspiration from Safe-RLHF [dai2023safe] and PPO [schulman2017proximal], the learning phase of ISA involves iteratively solving the min-max problem defined in Equation 3. Specifically, we alternate between updating the VLA model parameters, $\theta$, and the Lagrange multipliers, $\lambda$. The reward and cost functions at each time step $t$ are defined as follows. The reward $r_{t}$ is a function of the current state $s_{t}$ and the language instruction $l$:
@@
Notably, Equation 5 defines the instantaneous cost $c_{t}$ at a single timestep, while Equation 1 defines the cumulative cost by summing these instantaneous costs over a trajectory. Equation 1 builds upon Equation 5.
@@
where the objective functions $\mathcal{L}_{R}$ and $\mathcal{L}_{C}$ optimize a policy $\pi_{\theta}$ under safety constraints. Let $\mathcal{D}_{\text{task}}$ denote a dataset of task instructions. A task instruction $l$ is sampled from $\mathcal{D}_{\text{task}}$, $\tau=(s_{0},a_{0},s_{1},\dots)$ represents a trajectory, and $\tau\sim\pi_{\theta}$ denotes the trajectory distribution dependent on $\pi_{\theta}$: $s_{0}\sim\mu$, $a_{t}\sim\pi_{\theta}(\cdot|l,h_{t})$, $s_{t+1}\sim\mathbb{P}(\cdot|s_{t},a_{t})$. At each time step $t$, the policy considers a temporal context window defined by $h_{t}=\{(o_{t-n},a_{t-n}),(o_{t-n+1},a_{t-n+1}),\dots,(o_{t-1},a_{t-1}),o_{t}\}$, which contains the history of the past $n$ state-action pairs along with the current state $s_{t}$. The importance sampling ratio $\rho_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|l,h_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|l,h_{t})}$ measures the policy update magnitude relative to an old policy $\pi_{\theta_{\text{old}}}$. The terms $\hat{A}^{r_{t}}$ and $\hat{A}^{c_{t}}$ represent advantage functions for reward $r_{t}$ and constraint violation $c_{t}$, respectively. The $\operatorname{clip}(\rho_{t}(\theta),1-\epsilon,1+\epsilon)$ operator restricts $\rho_{t}(\theta)$ to $1-\epsilon,1+\epsilon]$, ensuring stable policy updates through proximal optimization. The combined loss $\mathcal{L}$ balances reward maximization and constraint satisfaction Lagrangian multiplier $\lambda$, where $\lambda\to 0$ prioritizes reward and $\lambda\to\infty$ enforces strict constraint adherence. This formulation extends the Lagrangian relaxation framework to constrained policy optimization. The method for updating the model parameters and Lagrange multipliers is as follows:
@@ C.2 Hyperparameters
Table 12: Hyper-parameters for training. We use AllenAct [weihs2020allenact] and OmniSafe
[ji2024omnisafe] as the training framework. Methods ISA FLaRe-Reward Shaping initial-lagrange-multiplier 0.001 N/A lagrange-multiplier-learning-rate 0.035 N/A total-rollouts 32 32 distributed-sampling-gpus 8 8 envs-per-device 4 4 actor-learning-rate 2.00E-5 2.00E-5 critic-learning-rate 2.00E-5 2.00E-5 actor-LR-scheduler-type constant constant critic-LR-scheduler-type constant constant iterations-per-update 1 1 update-repeats 4 4 clip-range-ratio 0.1 0.1 max-gradient-norm 0.5 0.5 discount-factor-$\gamma$ 0.99 0.99 gae-$\lambda$ 0.95 0.95 value-loss-weight 0.5 0.5 entropy-loss-weight 0.0 0.0 steps-per-ppo-update 128 128 transformer-encoder-layers 3 3 transformer-encoder-hidden-dims 512 512 transformer-encoder-heads 8 8 casual-transformer-decoder-layers 3 3 casual-transformer-decoder-hidden-dims 512 512 casual-transformer-decoder-heads 8 8
@@ Appendix D Details of Safety Constraints
[…]
@@
4:for pointpoint in $S$ do $\triangleright$ Obtain all points in the map
5:  if pointpoint is reachablereachable and $(point[0]-p[0])^{2}+(point[1]-p[1])^{2}\leq r$ then
@@
7:  else if pointpoint is unreachableunreachable and $(point[0]-p[0])^{2}+(point[1]-p[1])^{2}\leq r$ then
@@
[…]
@@
[…]
@@
[…]
3:for culsterculster in ClustersClusters do
@@
[…]
@@
[…]
@@
10:   $U\leftarrow U\cup\{o\}$ $\triangleright$ Object responded to minor force → unstable
@@
[…]
@@
1:Dangerous Collection $D$, Status Change Objects $S$ $\triangleright$ Gas, handsaw, grenade, arrow etc.
@@ E.2 Evaluation Tasks
Table 13: Details of evaluation tasks.
@@ E.3 Evaluation Models
We evaluated the safety and task performance of our method alongside state-of-the-art approaches. Our comparative experiments involved three types of method and eight models, encompassing both fair and unfair experimental setups. In the fair experiments, we evaluated two models, FLaRe and FLaRe Reward Shaping, which share the same imitation learning foundation model as our ISA but employ different reinforcement learning processes and are trained for no fewer steps than ISA. In unfair experiments, we used models trained exclusively with imitation learning, including SPOC-DINOv2, SPOC-SigLip-S and SPOC-SigLip-L. The first two models were pre-trained on the CHORES tasks [ehsani2024spoc], aligning with our foundation model, while the third was trained on the CHORES-L tasks using a larger imitation learning dataset than that used for our foundation model. Poliformer is a model trained from scratch using reinforcement learning and is only capable of performing the ObjNav task. Additionally, we incorporated two models equipped with privileged information, specifically visual bounding boxes for target objects. Following extensive evaluation and analysis, our method achieved state-of-the-art performance in both safety and task performance.
@@ Appendix F Related Work
Safety in Robotics. Safety in robotics has been a central focus of both the control and reinforcement learning communities [brunke2022safe], with the goal of ensuring robust safety guarantees and achieving generalization to previously unseen scenarios [aswani2013provably]. Traditional methods typically model and enforce safety constraints explicitly in analytical dynamic models, such as constrained motion planning [lozano2014constraint]. These constraints can include spatial limitations [saveriano2019learning], object pose restrictions and joint torque bounds [berenson2009manipulation], etc. However, these methods struggle with generalization to diverse scenarios [hewing2020learning]. In contrast, learning-based approaches typically rely less on prior knowledge, but their black-box nature makes it challenging to guarantee safety rigorously [koller2018learning]. Many previous works have explored the integration of control theory with reinforcement learning [dalal2018safe, thananjeyan2021recovery, marvi2021safe], focusing primarily on 1) learning dynamic models to predict unsafe consequences [hewing2020learning], 2) explicitly modeling safety in the objective function to encourage safe behaviors [kahn2017uncertainty], and 3) providing provable safety [luo2021learning]. Our work demonstrates that the paradigm of constrained learning can scale to large VLA models, leading to safety decisions that align with human values, which is highly relevant to 2).
@@ Appendix G Limitations and Future Work
Limitations. Despite the promising results, this work has several limitations. A primary limitation is the reliance on simulation for both training and evaluation. While prior work supports the feasibility of sim-to-real transfer for VLAs [ehsani2024spoc, hu2024flare], and simulation is indispensable for affordably collecting diverse safety-critical data, extensive validation on physical robotic platforms is a necessary next step.
Methodologically, our current framework employs several design choices that can be further refined. First, for trajectory-level violations, we assign cost credit to the final step of an unsafe sequence. While this approach avoids potential biases from hand-crafted reward shaping, exploring more advanced, heuristic-based credit assignment strategies is a promising direction to improve sample efficiency. Second, our safety constraints are binary and applied uniformly, rather than being explicitly linked to specific task instructions or weighted by severity. This uniform application is a simplification, as the notion of severity is often highly context-dependent.
Future Work. Building on these limitations, our future work will proceed in several key directions. The most immediate goal is to bridge the sim-to-real gap by validating and adapting the ISA framework on complex, real-world robotic platforms. This will involve tackling challenges like physical interaction dynamics and the irreversible consequences of failures.
To enhance the sophistication of our safety framework, we plan to move beyond the current constraint structure. A valuable next step is to leverage our framework’s extensibility to incorporate severity-weighted constraints, enabling more nuanced safety alignment tailored to specific applications and user preferences. We also aim to develop dynamic safety constraints that can adapt to changing environmental conditions and language-based instructions.
Furthermore, we plan to explore richer safety paradigms beyond expected cost minimization. This includes incorporating risk-sensitive metrics like Conditional Value at Risk (CVaR) to more effectively mitigate low-tail risks, and developing robust uncertainty estimation methods for real-time risk assessment, which could trigger more conservative policies when the model is uncertain. Ultimately, our vision is to develop a comprehensive, layered safety system that integrates algorithmic safeguards, adaptive mechanisms, and necessary physical safety measures to ensure robust and reliable deployment of embodied agents in the real world.